### v2

@@ 3 Random Attention Model with Unobserved References: Independent Random Attention
In this section I assume that the attention function in the RAM-UR model is of an independent random type in the sense that there exists an attention parameter $\gamma:X\rightarrow(0,1]$ such that the attention function takes the form: $\mu_{IRA}(D,A)=\prod_{x\in D\subseteq A}\gamma(x)\prod_{y\in A\setminus D}(1-(\gamma(y))$ for all $A\in\mathcal{X}$. Here attention privilege manifests in the form of alternatives with the attention parameter to be one, which is the key difference between this form and the attention function of the IRA form introduced by Manzini and Mariotti (2014) (hereafter referred to as the IRA model). The IRA restricts the attention parameter to be strictly between zero and one, we do not have such a restriction. This makes the IRA model to be nested in the RAM-UR-IRA.
@@
Imposing the above additional structure significantly changes how the probabilities can be defined in the RAM-UR-IRA model. Even though it is a special case of RAM-UR and nests [^8] By nesting the IRA (rational) I mean that a special case of RAM-UR-IRA will also be an IRA (rational) model. both the IRA as well as classical rationality, it is mathematically a very small perturbation of the IRA model: It allows $\gamma$ to take the value one for some alternatives. I provide a characterization of this model by using axioms similar to the characterizations of both the IRA and the RAM-UR and discuss the identification results.
@@ 3.1 Characterization
I characterize the RAM-UR-IRA using the following five axioms:
@@
This axiom is similar to the i-asymmetry axiom that characterizes the IRA (cf. Manzini and Mariotti (2014)). However, it constrains the denominator to be non-zero. Such a constraint was not needed in the IRA characterization, since all probabilities there are assumed to be positive. This axiom says that if removing an alternative $b$ changes the choice probability of another, say $a$ (given the positive denominator constraints), then removing $a$ would not change the choice probability of $b$.
R-Independence (R-IND): For all $A,B\in\mathcal{X}$ and $a,b\in A\cap B$ with $p(a,A)>0,p(a,B)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{p(a,B\setminus\{b\})}{p(a,B)}$.
For all $A,B\in\mathcal{X}$ and $b\in A\cap B$ with $p(a^{*},A)>0,p(a^{*},B)>0$: $\frac{p(a^{*},A\setminus\{b\})}{p(a^{*},A)}=\frac{p(a^{*},B\setminus\{b\})}{p(a^{*},B)}$.
@@
This axiom is similar in spirit to the EDA axiom used in the characterization of RAM-UR. However, we only need EDA*, which is restricted to checking it for the default alternative, as opposed to checking for all alternatives that are chosen with zero probability in a menu. Note again that EDA implies EDA*, which means that NT and EDA* together are weaker than EDA.
@@
This is the standard regularity axiom which states that an alternative’s choice probability weakly reduces in a larger menu. This property is in general not satisfied by the RAM-UR. If we consider the IRA model, the i-asymmetry and the i-independence axioms imply the regularity property. However, that is not the case with the other four axioms utilized to characterize the RAM-UR-IRA. The property is, thus, used as an axiom for the characterization. Note that this axiom implies that if we do not impose non-triviality, then any trivial alternative [^9] An alternative $x\in X$ here is called trivial if $p(x,\{x\})=0$. will not be chosen with a positive probability in any menu.
@@
Representation uniqueness means that the set of references, the underlying preferences, and the underlying attention parameters are uniquely identified. This uniqueness result is due to the additional structure imposed on the attention function, and the availability of data in menus with no reference points.
The proof of Theorem 2 essentially separates the menus among those with no references and ones with at least one reference. Using the axioms above and the Manzini and Mariotti (2014) result on the set of menus with no reference, it identifies the underlying attention parameter and partial preferences among non reference alternatives (precisely the part of the preferences that were unidentified in the RAM-UR). Since the RAM-UR-IRA is a special case of the RAM-UR, the rest of the identification is easily done similar to the general model. The proof shows that the fitting of the identified preferences through the non-reference menus and the preferences identified through the menus with references does indeed give us a strict total order. Finally, using the first two axioms, we obtain the required functional form.
@@ 4 Literature Overview
The idea that limited attention affects behavior and choice is very natural and well-documented in the literature. Simon’s (1956) model of satisficing choice is one of the earliest models that can be motivated by limited attention, here the heuristic is that the DM stops at the first alternative which is “good enough”. A DM might utilize such a heuristic because of lack of time or attention. Lleras et al. (2017) model choice overload which leads to worse outcomes in larger menus, the reason being that good options might fail to grab attention in larger sets due to an overload of available options. Masatlioglu et al. (2012) introduce the concept of the attention filter which incorporates the idea that removing an unconsidered alternative does not change the consideration set in a deterministic choice setting where the DM is attention constrained. Dean et al. (2017) is another paper that incorporate limited attention and also the effect of reference alternatives.
@@
The advent of stochastic choice as a fruitful direction of research to allow for the variability observed in choices called for modeling limited attention in such a framework. Manzini and Mariotti (2014) was the seminal paper that filled this void. Since then there has been plenty of literature combining both limited attention and stochastic choice. There are certainly lots of papers that are related closely with Manzini and Mariotti (2014), for instance, Brady and Rehbeck (2016)’s feasibility model is a generalization of the Manzini and Mariotti (2014), though their motivation is not limited attention, but limited feasibility and another one is Horan (2019) which looks at the IRA without a default. Cattaneo et al. (2020) introduces the general idea of a random attention model. Kovach and Suleymanov (2023) uses the idea of a RAM in the context of reference-dependent choice and modeled a family of attention functions, each indexed by the reference point, and looked at different special cases like the Constant Random Attention, Luce Random Attention and the Independent Random Attention. These authors discussed the nestedness properties among these special cases. They were not able to conduct the nestedness analysis with respect to the classical stochastic choice models because in their model the primitives were menu-reference pairs which put it outside the realm of stochastic choice models.
@@
The idea of reference-dependent choice has been the subject of extensive study as well. Masatlioglu and Ok (2005) and Masatlioglu and Ok (2014) are seminal papers in this area. These authors were one of the first to model status quo bias in choice, which is a widely observed behavior in experiments on individual choice [^10] Masatlioglu and Ok (2005) provides a few sources for such experiments.. Sagi (2006) imposes a no-regret condition on the reference dependent preferences and Tapkı (2007) allows for incomplete reference dependent preferences. Note that all mentioned works model the effect of the status quo on preferences. However, it is possible that the status quo might affect attention as well. Dean et al. (2017) looks at status quo bias and limited attention together, allowing for both reference-dependent preferences and attention. Kovach and Suleymanov (2023), as discussed before, constructs a model where references affect only the attention. Bossert and Sprumont (2009) provides another model where reference points don’t affect preferences. They model a deterministic choice function where the choice is consistent with not choosing a worse alternative than the reference. They characterize where there exists a linear order such that the choice from each menu-reference pair is at least as good as the reference point, not necessarily the maximum [^11] The best alternative according to the preference ordering.. The RAM-UR does have this non-deteriorating property, since references are attention privileged, so in any menu an alternative worse than the reference is never chosen.
@@
In the literature discussed above, a general theme is that the reference point is included as part of the primitive. It is natural to challenge the assumption of the observability of the reference point. Ok et al. (2015) and Tserenjigmid (2019) are two relevant papers here. However, in Tserenjigmid (2019) the references are not alternatives but a vector of attributes which depend upon the minimum of those attributes among the alternatives in any menu. It is thus, menu dependent as well. Kovach and Suleymanov (2023) is the first paper to study reference dependence in a stochastic setting, which makes my paper the first to study unobservable references in this framework.

### v3

@@ 4 Random Attention Model with Unobserved References: Independent Random Attention
In this section I assume that the attention function in the RAM-UR model is of an independent random type in the sense that there exists an attention parameter $\gamma:X\rightarrow(0,1]$ such that the attention function takes the form: $\mu_{IRA}(D,A)=\prod_{x\in D\subseteq A}\gamma(x)\prod_{y\in A\setminus D}(1-\gamma(y))$ for all $A\in\mathcal{X}$. Here attention privilege manifests in the form of alternatives with the attention parameter to be one, which is the key difference between this form and the attention function of the IRA form introduced by Manzini and Mariotti (2014) (hereafter referred to as the IRA model). The IRA restricts the attention parameter to be strictly between zero and one while we do not impose such a restriction. This makes the IRA model to be nested in the RAM-UR-IRA.
@@
Imposing the above additional structure significantly changes how the probabilities can be defined in the RAM-UR-IRA model. Even though it is a special case of RAM-UR and nests [^11] By nesting the IRA (rational) I mean that a special case of RAM-UR-IRA: the one where $E=\phi$ will also be an IRA (rational) model. both the IRA as well as classical rationality, it is mathematically a small departure from the IRA model: It allows $\gamma$ to take value one for some alternatives. In what follows, I provide a characterization of this model by using axioms similar to the characterizations of both the IRA and the RAM-UR and discuss the identification results.
@@ 4.1 Characterization
I characterize the RAM-UR-IRA using the following five independent axioms:
@@
This axiom is similar to the i-asymmetry axiom that characterizes the IRA (cf. Manzini and Mariotti (2014)). However, it restricts the denominator to be non-zero. Such a restriction was not needed in the IRA characterization, since all probabilities are assumed to be positive in that model. The axiom says that if removing an alternative $b$ changes the choice probability of another, say $a$ (given the positive denominator constraints), then removing $a$ would not change the choice probability of $b$.
R-Independence (R-IND): (1) For all $A,B\in\mathcal{X}$ and $a,b\in A\cap B$ with $p(a,A)>0,p(a,B)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{p(a,B\setminus\{b\})}{p(a,B)}$.
(2) For all $A,B\in\mathcal{X}$ and $b\in A\cap B$ with $p(a^{*},A)>0,p(a^{*},B)>0$: $\frac{p(a^{*},A\setminus\{b\})}{p(a^{*},A)}=\frac{p(a^{*},B\setminus\{b\})}{p(a^{*},B)}$.
@@
This axiom is similar in spirit to the EDA axiom used in the characterization of RAM-UR. However, we only need a check for the default alternative as opposed to checking for all alternatives that are chosen with zero probability in a menu. Note again that EDA implies EDA*, which means that NT and EDA* together are weaker than EDA.
@@
This is the standard regularity axiom which states that an alternative’s choice probability weakly reduces in a larger menu. This property is in general not satisfied by the RAM-UR. If we consider the IRA model, the i-asymmetry and the i-independence axioms imply the regularity property. However, that is not the case with the other four axioms utilized to characterize the RAM-UR-IRA. Thus, the property is used as an axiom for the characterization. In the absence of non-triviality, this axiom implies that any trivial alternative [^12] An alternative $x\in X$ here is called trivial if $p(x,\{x\})=0$. will not be chosen with a positive probability in any menu.
@@
Representation uniqueness means that the set of references, the underlying preferences, and the underlying attention parameters are uniquely identified. This uniqueness result is due to the additional structure imposed on the attention function and to the availability of data in menus with no reference points.
The proof of Theorem 2 essentially partitions the menus into those with no references and ones with at least one reference. Using the axioms above and the Manzini and Mariotti (2014) result on the set of menus with no reference, it identifies the underlying attention parameter and partial preferences among non reference alternatives (precisely the part of the preferences that were unidentified in the RAM-UR). Since the RAM-UR-IRA is a special case of the RAM-UR, the rest of the identification is easily done similar to the general model. The proof shows that the fitting of the identified preferences through the non-reference menus and the preferences identified through the menus with references does indeed give us a strict total order. Finally, using the first two axioms, we obtain the required functional form.
@@ 5 Literature Overview
The idea that limited attention affects behavior and choice is well accepted and widely documented in the literature. Simon’s (1956) model of satisficing choice is one of the earliest models that can be motivated by limited attention. In this model, the heuristic is that the DM stops at the first alternative which is “good enough”. A DM might utilize such a heuristic because of lack of time or attention. Lleras et al. (2017) model “Choice Overload” which leads to worse outcomes in larger menus, the reason being that good options might fail to grab attention in larger sets due to an overload of available options. Masatlioglu et al. (2012) introduce the concept of the attention filter which incorporates the idea that removing an unconsidered alternative does not change the consideration set in a deterministic choice setting where the DM is attention constrained. Dean et al. (2017) is another paper that incorporates limited attention and also the effect of reference alternatives.
@@
Models of stochastic choice ventured into modelling the variability observed in choices due to limited attention. Manzini and Mariotti (2014) was the seminal paper that filled this void. Since then there has been a substantial body of literature combining both limited attention and stochastic choice. A plethora of papers are closely related to Manzini and Mariotti (2014), for instance, Brady and Rehbeck (2016)’s feasibility model is a generalization of the Manzini and Mariotti (2014), though their motivation is not limited attention, but limited feasibility. Another one is Horan (2019) which looks at the IRA without a default. Cattaneo et al. (2020) introduces the general idea of a random attention model. Kovach and Suleymanov (2023) uses the idea of a RAM in the context of reference-dependent choice and models a family of attention functions, each indexed by the reference point, and looked at different special cases like the Constant Random Attention, Luce Random Attention and the Independent Random Attention. They discussed the nestedness properties among these special cases. They were unable to conduct the nestedness analysis with respect to the classical stochastic choice models because in their model, the primitives were menu-reference pairs, which placed it outside the realm of stochastic choice models.
@@
The idea of reference-dependent choice has also been the focus of several papers. Masatlioglu and Ok (2005) and Masatlioglu and Ok (2014) are seminal papers in this area. These authors were one of the first to model status quo bias in choice, which is a widely observed behavior in experiments on individual choice [^13] Masatlioglu and Ok (2005) provides a few references for such experiments.. Sagi (2006) imposes a no-regret condition on the reference dependent preferences and Tapkı (2007) allows for incomplete reference dependent preferences. Note that all mentioned works model the effect of the status quo on preferences. However, it is possible that the status quo might affect attention as well. Dean et al. (2017) looks at status quo bias and limited attention together, allowing for both reference-dependent preferences and attention. Kovach and Suleymanov (2023), as discussed before, constructs a model where references affect only the attention. Bossert and Sprumont (2009) provides another model where reference points don’t affect preferences. They model a deterministic choice function where the choice is consistent with not choosing a worse alternative than the reference. They study the existence of a linear order such that the choice from each menu-reference pair is at least as good as the reference point, though not necessarily the maximum [^14] The best alternative according to the preference ordering.. The RAM-UR does have this non-deteriorating property since references are attention privileged, so in any menu, an alternative worse than the reference is never chosen.
@@
In the literature discussed above, a general theme is that the primitives include the reference point. It is natural to challenge the assumption of the observability of the reference point. Ok et al. (2015) and Tserenjigmid (2019) are two relevant papers here. However, in Tserenjigmid (2019) the references are not alternatives but a vector of attributes which depend upon the minimum of those attributes among the alternatives in any menu. It is thus, menu dependent as well. Kovach and Suleymanov (2023) is the first paper to study reference dependence in a stochastic setting, which makes the current work the first to study unobservable references in this setup.